A double purpose tank 18 ft wide, 24 ft long and 4 ft depth is filled with water. What is the weight of water in long tons? in 2 decimal places


1
Expert's answer
2021-08-16T08:32:58-0400

"m = ρ \times V"

ρ = density of water

ρ = 62.4 lb/ft3

"V = 18 \times 24 \times 4 = 1728 \;Ft^3 \\

m = 1728 \times 62.4 = 107827.2 \;kg \\

W = mg \\

W= 107827.2 \times 32.17 = 3468801.02 \;N"

Answer: 3468801.02 N